### v1

@@ IV Discussions
The $t+t$ RWA results of the ground state, the first excited state, and the $t+t$ resonance states are shown in Fig. 3 - Fig. 7. The RWA has considerable amplitudes for the ground and the first excited states when the inter-cluster distance $a$ is small and decreases rapidly for larger $a$. In contrast, the $t+t$ resonance states are featured with extended amplitudes. The numbers of nodes in their RWA curves increase as the energies are excited higher. As shown in Eq. (10), the RWA is defined as the overlap amplitude between the wave function of ${}^{6}\mathrm{He}$ and the coupling wave functions of two $t$ clusters. As a result, the absolute value of RWA directly reflects the probability of cluster formation at a given distance $a$, while the number of nodes corresponds to the excitation of inter-cluster motion.
The current calculation provides a robust and consistent description with the coupling configurations of $\alpha+n+n$ and $t+t$, not only characterize the system at the lowest energy levels but also offer an accurate description of the $t+t$ states above the threshold, highlighting their relevance in the broader spectrum. To delve deeper into the properties of these states, we have calculated the monopole transition strengths within the isoscalar monopole (ISM). Specifically, the monopole transition from the ground state $0_{1}^{+}$ to the third excited state $0_{3}^{+}$ has a strength of $B(ISM;0_{1}^{+}\to 0_{3}^{+})=2.8$ $\text{fm}^{4}$, while the transition from $0_{1}^{+}$ to the fourth excited state $0_{4}^{+}$ is slightly stronger, with $B(ISM;0_{1}^{+}\to 0_{4}^{+})=4.9$ $\text{fm}^{4}$. In contrast, the transition between the third and fourth excited states, $0_{3}^{+}\to 0_{4}^{+}$, shows a significantly larger transition strength of $B(ISM;0_{3}^{+}\to 0_{4}^{+})=48.6$ $\text{fm}^{4}$. These values offer important insights into the transition probabilities between the low-lying and excited states, providing a quantitative measure of the nuclear structure through isospin transitions. The relatively large value of $B(ISM;0_{3}^{+}\to 0_{4}^{+})$ compared to the others suggests a strong transition between the third and fourth excited states, which points to a dominant role for the $t+t$ configuration at higher energy levels. In contrast, the lower values of $B(ISM;0_{1}^{+}\to 0_{3}^{+})$ and $B(ISM;0_{1}^{+}\to 0_{4}^{+})$ indicate that the ground state is primarily governed by the $\alpha+n+n$ component.
@@
The GCM calculations indicate that the $t+t$ cluster resonant states above the threshold exhibit a rich variety of quantum state characteristics. This richness arises primarily from the fermionic nature of the $t$ cluster, resulting in significantly more quantum energy levels for the $t+t$ resonance compared to the $\alpha+\alpha$ resonance in 8Be. Due to the lack of precise measurements and the inability to extract the spin-parity of the observed resonant state, it is impossible to confirm which quantum state the observed energy level corresponds to. Consequently, effectively comparing the energy levels measured in different experiments becomes challenging. Furthermore, multiple quantum states are permissible for $t+t$ resonance near the threshold, with experimental measurements concentrated in this region, as indicated by the black-colored columns in FIG.  2. Therefore, accurately measuring the $t+t$ resonance state in experiments is essential.
@@ V Conclusions
The 6He isotope is a neutron-rich nucleus that exhibits above-threshold $t$+$t$ cluster states, providing valuable insights into the behaviour of loosely bound nuclear systems and their formation mechanisms. The discovery of multiple resonance states is particularly significant as it contrasts with the traditional $\alpha$ cluster structures observed in other light nuclei, highlighting the unique fermionic properties of the triton cluster. The excitation energy spectrum above the $t$+$t$ threshold in 6He is measured during the decay of 9Li on a 208Pb target at 32.7 MeV/nucleon. The resonant states are reconstructed using the invariant mass method and reveal peaks at $13.9\pm 0.3$, $15.0\pm 0.3$, $17.016\pm 0.002$ and $19.4\pm 0.6$ MeV, the latter two peaks being reported for the first time. These results are compared with previous observations, although inconsistencies remain. Microscopic cluster model calculations, including the coupling between $t+t$ and $\alpha+n+n$ configurations, are carried out to explore the $t+t$ resonance states in 6He. The theoretical energy spectra are in good agreement with the experimental results. To further investigate the clustering structure of the theoretical states, we calculate the RWA of the $t+t$ channels. As expected, the identified $t+t$ resonance states show significant amplitudes in the $t+t$ RWA, reinforcing their clustering nature. This study highlights the complexity of the energy level discrepancies in the excited states of 6He and provides accurate theoretical predictions. These results underline the need for more accurate and comprehensive experimental studies to further explore the structure of the $t+t$ resonance in 6He. A future experimental setup should feature high detection efficiency, high statistics, and high precision to accurately determine resonance states with spin-parity. Angular correlation analysis is a valid method for experimentally identifying spin-parity for the sequential breakup reaction described as $a(A,B^{*}\rightarrow c+C)b$  [64, 65].
@@ Acknowledgements
This work is supported by the National Key R&\&&D Program of China (2023YFA1606404,2023YFA1606701,2022YFA1602402). This work is supported in part by the National Natural Science Foundation of China under contract Nos. 12175042, 11890710, 11890714, 12047514, 12147101, 12475138, 11875297, and 12347106, the Guangdong Major Project of Basic and Applied Basic Research No. 2020B0301030008, the STCSM under Grant No. 23590780100, and the Natural Science Foundation of Shanghai under Grant No. 23JC1400200.

### v2

@@ IV Discussions
Fig. 4 shows the $t+t$ RWA results for the ground state, the first excited state, and two positive-parity $t+t$ resonances, whereas Fig. 5 shows the results for negative-parity $t+t$ resonances. The RWA has considerable amplitudes for the ground and the first excited states when the inter-cluster distance $a$ is small and decreases rapidly for larger $a$. In contrast, the $t+t$ resonant states are featured with extended amplitudes. As shown in Eq. (10), the RWA is defined as the overlap amplitude between the wave function of $^{6}\mathrm{He}$ and the coupling wave functions of two $t$ clusters. As a result, the absolute value of RWA reflects the probability of cluster formation at a given distance $a$.
In the GCM analysis of the ${{}^{3}\mathrm{He}}+{{}^{3}\mathrm{He}}$ scattering system [51], the inter-cluster orbital angular momentum was identified as a critical quantity determining state characteristics, due to spin-coupling effects inherent to fermion systems. In the present study, through RWA results, we identify prominent angular momentum channels across various states. For positive-parity states, the triton clusters form spin-singlet states with $s$-wave and $d$-wave relative motion, respectively. Conversely, the negative-parity states exhibit spin-triplet states, where $p$-wave dominates in the $1^{-}$ and the first $2^{-}$ states, and $f$-wave dominates in the $3^{-}$ and the second $2^{-}$ states.
The current calculation provides a robust and consistent description with the coupling configurations of $\alpha+n+n$ and $t+t$, not only characterize the system at the lowest energy levels but also offer an accurate description of the $t+t$ states above the threshold, highlighting their relevance in the broader spectrum.
@@
The GCM calculations indicate that the $t+t$ cluster resonant states above the threshold exhibit a rich variety of quantum state characteristics. This richness arises primarily from the fermionic nature of the $t$ cluster, resulting in significantly more quantum energy levels for the $t+t$ resonance compared to the $\alpha+\alpha$ resonance in 8Be. Due to the lack of precise measurements and the inability to extract the spin-parity of the observed resonant state, it is impossible to confirm which quantum state the observed energy level corresponds to. Consequently, effectively comparing the energy levels measured in different experiments becomes challenging. Furthermore, multiple quantum states are permissible for $t+t$ resonance near the threshold, with experimental measurements concentrated in this region, as indicated by the black-colored columns in FIG.  3. Therefore, accurately measuring the $t+t$ resonant state in experiments is essential.
@@ V Conclusions
The 6He isotope is a neutron-rich nucleus that exhibits above-threshold tt+$t$ cluster states, providing valuable insights into the behaviour of loosely bound nuclear systems and their formation mechanisms. The discovery of multiple resonant states is particularly significant as it contrasts with the traditional $\alpha$ cluster structures observed in other light nuclei, highlighting the unique fermionic properties of the triton cluster. The excitation energy spectrum above the tt+$t$ threshold in 6He is measured during the decay of 9Li on a 208Pb target at 32.7 MeV/nucleon. The resonant states are reconstructed using the invariant mass method and reveal peaks at $13.9\pm 0.3$ and $15.0\pm 0.3$ MeV. These results are compared with previous observations, although inconsistencies remain. Microscopic cluster model calculations, including the coupling between $t+t$ and $\alpha+n+n$ configurations, are carried out to explore the $t+t$ resonant states in 6He. The theoretical energy spectra are in good agreement with the experimental results. To further investigate the clustering structure of the theoretical states, we calculate the RWA of the $t+t$ channels. As expected, the identified $t+t$ resonant states show significant amplitudes in the $t+t$ RWA, reinforcing their clustering nature. This study highlights the complexity of the energy level discrepancies in the excited states of 6He and provides accurate theoretical predictions. These results underline the need for more accurate and comprehensive experimental studies to further explore the structure of the $t+t$ resonance in 6He. A future experimental setup should feature high detection efficiency, high statistics, and high precision to accurately determine resonant states with spin-parity. Angular correlation analysis is a valid method for experimentally identifying spin-parity for the sequential breakup reaction described as $a(A,B^{*}\rightarrow c+C)b$  [71, 72].
@@ Acknowledgements
This work is supported by the National Key R&\&D Program of China (2023YFA1606404,2023YFA1606701,2022YFA1602402). This work is supported in part by the National Natural Science Foundation of China under contract Nos. 12175042, 11890710, 11890714, 12047514, 12147101, 12475138, 11875297, and 12347106, the Guangdong Major Project of Basic and Applied Basic Research No. 2020B0301030008, the STCSM under Grant No. 23590780100, and the Natural Science Foundation of Shanghai under Grant No. 23JC1400200.